OpenStudy (anonymous):

could someone just give me the steps for solving this prob? Solve the equation by using factoring: x2 -14x + 49 =0

OpenStudy (anonymous):

start by looking at the factors of 49

OpenStudy (anonymous):

7*7

OpenStudy (anonymous):

break the mid term x^2-7x-7x+49 now factor the first two terms and last two terms .

OpenStudy (anonymous):

Factorise the quadratic so that you get something of the form \((x\pm a)(x\pm b)\). With a quadratic of the form \(ax^2+bx+c\), you generally should look to find products of \(c\) that sum to \(b\).

OpenStudy (anonymous):

i havent learned quads yet

OpenStudy (anonymous):

so x^2-7x=0 -7x+49=0 and i try to solve for x in each one?

OpenStudy (anonymous):

no x^2-7x-7x+49=0 x*(x-7)-7(x-7)=0 (x-7)(x-7)=0

OpenStudy (anonymous):

um these are my answers 14 7 2 no solution

OpenStudy (anonymous):

the answer is 7?

OpenStudy (anonymous):

Factorising is effectively finding the roots, the same as using the quadratic formula (although factorising isn't always possible if the roots aren't integer). A quadratic is any equation of the form \(ax^2+bx+c\) - in your case, a = 1, b = -14 and c = 49. Find the products of \(c\): [1,49], [-1,-49], [7,7], [-7,-7] From the products, find which pair sum to \(b\): -7 + -7 = -14. Then you have the factorised form: \((x-7)(x-7)\). So you know the roots will be 7, as this will result in at least one (in this case, both) terms equating to zero.

OpenStudy (anonymous):

yes its 7

OpenStudy (anonymous):

myhomework says to find it by factoring

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

I've just told you how to factor it.